OpenStudy (anonymous):

How can I solve equation x^2 - 4y - 1 = 0, where x and y are natural numbers.

OpenStudy (ja1):

by natural numbers you refer to 1?

OpenStudy (anonymous):

By natural numbers I mean the set of all positive integers.

OpenStudy (anonymous):

I found some solutions. But is there any general formula for x and y?

OpenStudy (anonymous):

For example x = 3, y = 2 is a solution.

OpenStudy (anonymous):

x^2-4y-1=0 4y=x^2-1 x>0,y>0.x&y are natural numbers \[y=\frac{ x ^{2}-1 }{4 }\] Least positive value of x which satisfies the above condition is x=3 for x=3,y=(3*3-1)/4=2 add multiple of 4 in x=3 ,we get infinite values of y